OpenStudy (anonymous):

SOMEONE PLEASE HELP!!! The perimeter of a rectangle is 96 ft. Find the dimensions of the rectangle if the ratio of the length to the width is 7: 5. (Points : 1) 24 ft by 24 ft 28 ft by 20 ft 22 ft by 26 ft 30 ft by 18 ft

OpenStudy (anonymous):

will give medal

OpenStudy (perl):

in a rectangle you have two sides of 'length' and two sides of 'width'

OpenStudy (anonymous):

yes ok ok

OpenStudy (perl):

P = 2L + 2W

OpenStudy (perl):

you are given P = 96 and L : W = 7:5. Which means L / W = 7/5 , you can cross multiply

OpenStudy (anonymous):

5l=7w?

OpenStudy (perl):

correct

OpenStudy (perl):

now solve for L or W

OpenStudy (perl):

your choice

OpenStudy (anonymous):

may i ask, how do i solve?

OpenStudy (anonymous):

@perl

OpenStudy (perl):

so far you have the equations 96 = 2L + 2W 5L = 7W you can solve this system by substitution

OpenStudy (anonymous):

ohhhh ok

OpenStudy (anonymous):

@kamibug

OpenStudy (anonymous):

i was jk...warrior just wanted me to put it on >.>

OpenStudy (kamibug):

Haha, that's alright. xD